OpenStudy (anonymous):

Will medal and fan Find the maximum or minimum of the following quadratic function: y = -2x2 + 4x - 1. A. -1 B. 0 C. 1 D. 4

OpenStudy (anonymous):

I assume you mean -2x^2?

OpenStudy (anonymous):

y= -2x^2+4x-1

OpenStudy (anonymous):

Are you familiar with the quadratic formula?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

What you can do here is use the first bit of the quadratic formula (-b/2a) which will find you the x value of the vertex. Then you simply plug that x value back into your original equation to find the y value of the max / min. Since this function starts with -2, you know that it opens downwards so you will be finding a maximum.

OpenStudy (anonymous):

(-b/2a) or (-4/4) = (-1). This is the x value of the maximum. Plugging this back into the equation... -2(-1)^2 + 4(-1) - 1 = -2(1) - 4 - 1 = -2 - 4 - 1 or -7. Therefore we know the y value of the maximum is at -7. Seeing as -7 isn't one of your answers, I'm going to assume that the question simply wants you to find the x value of the maximum - that is, only do the first bit of all of this. :) Hope I helped!

OpenStudy (anonymous):

so the answer would be -1?

OpenStudy (anonymous):

Yes, it looks as though the answer would be -1. The question is slightly ambiguous in its wording, but since this function has a maximum at (-1,-7), -1 is the only coordinate that has a matching answer.

OpenStudy (anonymous):

Thank you for helping me